Please help me solve by completing the square, and if you can show me that would be great!
4. Solve by completing the square:
x2 + 6x + 1 = 0

Answer by stanbon(75887)

Solve by completing the square:
x^2 + 6x + 1 = 0
---------
x^2+6x+? = -1+?
x^2+6x + [(1/2)6]^2 = -1 + [(1/2)6]^2
x^2+6x+9 = -1+9
Factor:
(x+3)^2 = 8
Take the square root of both sides to get:
x+3 = 2sqrt2 or x+3 = -2sqrt2
x = -3+2sqrt2 or x = -3-2sqrt2
============
Cheers,
Stan H.